OpenStudy (anonymous):

Find (f*g)(x) for the indicated functions: f(x)=4x-6, g(x)= 5/x+6

OpenStudy (anonymous):

\[f\circ g(x)\]?

OpenStudy (anonymous):

yes, sorry I don't know how to make the little circle thing

OpenStudy (anonymous):

first write \[f(g(x))\] then replace g by what it is \[f(\frac{5}{x+6})\] then replace the x in f(x) by the new input to get \[4\times \frac{5}{x+6}-6\]

OpenStudy (anonymous):

and you can write \[f\circ g(x)=\frac{20}{x+6}-6\] but nothing cancels so you should be done at this step

OpenStudy (anonymous):

hmm, it says that isn't right

OpenStudy (anonymous):

well maybe they want to you actually subtract

OpenStudy (anonymous):

\[\frac{20}{x+6}-6=\frac{20}{x+6}-\frac{6(x+6)}{x+6}\] \[=\frac{20-(6x+36)}{x+6}\] \[=\frac{-6x-16}{x+6}\] but that seems like a lot of extra work to me

OpenStudy (anonymous):

that's it! thank you :-)